OpenStudy (anonymous):

Please help with the following question: Find rational numbers a and b such that: (7+5√2) ^ 1/3 = a+b√2 Thank you

hartnn (hartnn):

my first thoughts would be to cube both sides to eliminate the 1/3 exponent

hartnn (hartnn):

have u tried that ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

so, in which step are you stuck ?

OpenStudy (anonymous):

i get a multiple of complex equations

hartnn (hartnn):

did u get a(a^2+6b^2) =7 and b(2b^2+3a^2) =5 bdw, that ^ clearly suggests a=1,b=1 as one solution

OpenStudy (anonymous):

yes i got that

hartnn (hartnn):

you can just stop that because thats the only real solution!

hartnn (hartnn):

stop there***

hartnn (hartnn):

oh, and since you are new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{ }}\color{red}{\text{!}}\color{orange}{\text{!}}\color{#e6e600}{\text{}}\end{array} \]

OpenStudy (anonymous):

thanks